A 33 kg child sits in a 55 kg boat at rest, and throws a 4.1 kg package horizontally from the boat at a speed of 7 m/s. What is the velocity of the boat after this? (hint: the boy and the boat move together so consider that to be one object, the package is the other object)
1
Expert's answer
2018-01-19T15:59:07-0500
From the law of conservation of momentum
(m_1+m_2 )v=m_3 u
So
v=m_3/(m_1+m_2 ) u
v=4.1/(33+55)×7=0.33 m/s.